Barn painting. Melanie can paint a certain barn by herself
in x days. Her helper Melissa can paint the same barn by
herself in 2x days. Write a rational expression for the fraction
of the barn that they complete in one day by working together. Evaluate the expression for x  5.
: Barn painting. Melanie can paint a certain barn by herself
in x days. Her helper Melissa can paint the same barn by
herself in 2x days. Write a rational expression for the fraction
of the barn that they complete in one day by working together. Evaluate the expression for x  5.

Answer by ankor@dixie-net.com(4684)

Barn painting. Melanie can paint a certain barn by herself
in x days. Her helper Melissa can paint the same barn by
herself in 2x days. Write a rational expression for the fraction
of the barn that they complete in one day by working together
:
Let the completed job = 1
:
1/x + 1/((2x)) = (2 + 1)/(2x) = 3/(2x)
:
Evaluate the expression for x = 5
:
3/((2*5)) = 3/10 of the barn painted in one day
: